OpenStudy (anonymous):

In New York, the tax on a property assessed at $470,000 is $6,580. If tax rates are proportional in this city, how much would the tax be on a property assessed at $480,000? Choose the proportion you would use to solve the problem. How do I know how to set up the proportion like what do I need to do? Please!

OpenStudy (ybarrap):

6,580 = rate*470,000 so rate = 6,580/470,000, this is the proportion Then tax on 480,000 would be rate*480,000

OpenStudy (anonymous):

So the proportion would be.. 6,580/470,000=x/480,000